OpenStudy (anonymous):

A 5-foot ladder is leaning against a 4-foot wall. How far must the bottom of the ladder be from the base of the wall so that the top of the ladder rests on the top of the wall?

OpenStudy (sleepyjess):

 use the pythagorean theorem to figure this out

OpenStudy (sleepyjess):

\[a^2 +b^2=c^2\]

OpenStudy (sleepyjess):



OpenStudy (sleepyjess):

\[a^2=16\] \[b^2=?\] \[c^2=25\] \[16+x=25\] Once you get that take the square root of x and that is your answer.